# Optimal Rank-Metric Codes with Rank-Locality from Drinfeld Modules

### v1

@@ Abstract.
We introduce a new technique to construct rank-metric codes using the arithmetic theory of Drinfeld modules over global fields, and Dirichlet Theorem on polynomial arithmetic progressions. Using our methods, we obtain a new infinite family of optimal rank-metric codes with rank-locality, i.e. every code in our family achieves the information theoretical bound for rank-metric codes with rank-locality.
@@ 1. Introduction
Let $q$ be a prime power and $m$ a positive integer. Rank-metric codes are an important family of error correcting codes that are used for reliable communication over noisy channels [1, 5, 12, 18, 23] and for cryptography [7, 10, 21]. They have been widely studied in the literature due to their interesting algebraic combinatorial properties that allow to encode information in a way that the obtained codeword are resilient to rank errors, i.e. the noise is modeled as a sum of a low rank matrix.
@@
Rank-metric codes with rank-locality are an important subclass of rank-metric codes that allow efficient corrections of criss-cross failures in data centers (for additional detail on this see for example [14, Example 1]). Various constructions with flexible parameters are available for locally recoverable codes over the Hamming metric [4, 3, 6, 8, 11, 15, 16, 17] but only one for the rank-metric codes [14] (with very restrictive parameters). The difficulty that arises in the rank-metric context is that classical constructions fail due to the non-commutative nature of the message space, which naturally consists of $q$-linearized polynomials with coefficients in ${\mathbb{F}}_{q^{m}}$.
This paper focuses on general constructions of optimal rank-metric codes with rank-locality, i.e. codes whose dimension achieves the information theoretical limits imposed by the requiring locality and minimum distance. Our methods leverage powerful tools from the theory of Drinfeld modules. The shape of our message space is drawn by the decomposition of Anderson motives as free modules over ${\mathbb{F}}_{q^{m}}[T]$ (using the natural Drinfeld action). The evaluation points that provide the encoding are picked by carefully selecting an $h\in{\mathbb{F}}_{q}[T]$ and a Drinfeld module $\phi$ so that the $h$-torsion is defined of ${\mathbb{F}}_{q^{m}}$. Differently from [22] (where the authors get classical locality, as explained in [14, Section 3]), we actually get proper rank-locality, as required in [14]. Guaranteeing the existence of such $h$, relies on an effective version of Dirichlet Theorem on polynomial arithmetic progression, which we derive from the $L$-series technique in [20, Chapter 4].
For the rest of this introduction we describe our main information theoretical result. The reader who is not familiar with the rank-metric code material should refer to section 2.3 before diving into the coding theory result below. Our main result is the following.
Theorem 1.1.
$2\leq\delta<R$, $s+1\leq\ell<q$, and $q$ is coprime to $R$. Then, there exists a rank-metric code in ${\mathbb{F}}_{q}^{m\times\ell R}$, rank-locality $(r,\delta)$, $R=r+\delta-1$, and ${\mathbb{F}}_{q}$-dimension equal to $m(s+1)r$. Moreover, this code has rank distance $\ell R-Rs-r+1$, thus it is optimal.
Notice that most of these condition are non restrictive, but simply guarantee that there is enough space in ${\mathbb{F}}_{q}^{m\times\ell R}$ to build a optimal code with rank-locality having non-trivial parameters. For example, the code lives in a space of dimension $m\ell R$ and its ${\mathbb{F}}_{q}$-dimension will be $m(s+1)r$, so it is expected to have the inequality $s+1\leq\ell$, or for example $\delta\geq 2$ is necessary to have rank-locality and $\delta\leq R-1$ simply requires that the distance of the local code (i.e. the code in ${\mathbb{F}}_{q}^{m\times R}$ where one punctures at all columns except the ones in the locality set) makes sense.
@@ 2.1. Drinfeld Modules
Throughout the article, let $q$ be a prime power and ${\mathbb{F}}_{q}$ the finite field with $q$ elements. We denote by ${\mathbb{F}}_{q}[T]$ the ring of polynomials over ${\mathbb{F}}_{q}$ and by ${\mathbb{F}}_{q}(T)$ the rational function field in the variable $T$. A function field $F$ over ${\mathbb{F}}_{q}$ is denoted by $F/{\mathbb{F}}_{q}$, or simply $F$ whenever the constant field is understood. Let $P$ be a place of $F$. We denote by $\mathcal{O}_{P}$ the valuation ring at $P$. We will also use other standard notation for function fields, for which we refer to [24].
@@
In this section, we provide some background about the theory of Drinfeld modules. For a comprehensive treatise on the subject, refer to [19]. Throughout this section we assume $K\supseteq{\mathbb{F}}_{q}$ is a field and $a\in{\mathbb{F}}_{q}[T]$.
@@
$a\longmapsto$ $\phi_{a}=\gamma(a)+\sum^{n}_{i=1}h_{i}(a)\tau^{i}$
where $h_{n}(a)\neq 0$ and $n=\operatorname{deg}_{T}(a)\cdot R$.
The morphism $\gamma$ is called the structure morphism of $\phi$. Observe that for any $a=\sum_{i=0}^{n}a_{i}T^{i}$, we have $\phi_{a}=\sum_{i=0}^{n}a_{i}\phi_{T}^{i}$. In other words, $\phi_{T}$ determines $\phi$ entirely. The Carlitz module is the Drinfeld module of rank $1$ and defined by $\phi_{T}=\gamma(T)+\tau$ [19, Definition 3.2.5]. Moreover, $K\{\tau\}$ is related to the ring of $q$-linearized polynomials $K\langle x\rangle$ via the following isomorphism
@@
We say that $\iota(\phi_{a})$ is the $a$-division polynomial of $\phi$ and denote it by $\phi_{a}(x)$.
@@
It turns out that the set of $a$-torsion points has a more general structure than just an ${\mathbb{F}}_{q}$-vector space.
@@
For $K={\mathbb{F}}_{q}(T)$, we define good reduction at a prime ideal $\mathfrak{p}=(P)$ of ${\mathbb{F}}_{q}[T]$. But first, we need to recall the definition of non-Archimedean valuation on ${\mathbb{F}}_{q}(T)$.
@@
Let $P$ be a monic irreducible polynomial in ${\mathbb{F}}_{q}[T]$ and $\mathfrak{p}$ the ideal generated by P. For any nonzero $h\in{\mathbb{F}}_{q}(T)$, we can write $h=P^{n}\cdot\frac{f}{g}$ for some $n\in\mathbb{Z}$ with $f,g\in{\mathbb{F}}_{q}[T]$, $\operatorname{gcd}(f,g)=1$ and $g\neq 0$ such that $P\nmid fg$. The discrete valuation $v_{\mathfrak{p}}:{\mathbb{F}}_{q}(T)\rightarrow\mathbb{Q}\cup\{\infty\}$ is defined as
$v_{\mathfrak{p}}(h)=\begin{cases}n&\text{ if }h\neq 0\\ \infty&\text{ if }h=0\\ \end{cases}$
@@
[19, Remark 7.1.18] Let $\mathfrak{p}$ be a prime ideal in ${\mathbb{F}}_{q}[T]$ and ${\mathbb{F}}_{\mathfrak{p}}$ be the residue field at $\mathfrak{p}$. Let $\phi:{\mathbb{F}}_{q}[T]\rightarrow{\mathbb{F}}_{q}(T)\left\{\tau\right\}$ be a Drinfeld module of rank $R$. Let $\phi_{T}=\sum_{i=0}^{R}a_{i}\tau^{i}$. We say that $\phi$ has a good reduction at $\mathfrak{p}$ if it satisfies the following conditions:
  • $v_{\mathfrak{p}}(a_{R})=0$,
  • $v_{\mathfrak{p}}(a_{i})\geq 0$ for $10\leq i\leq R-$ 1.
If these conditions are verified, then $\overline{\phi}:{\mathbb{F}}_{q}[T]\rightarrow{\mathbb{F}}_{\mathfrak{p}}\{\tau\}$ is given by reducing the coefficients of $\phi_{T}$ at $\mathfrak{p}$ and naturally extending the homomorphism using the algebra structure of ${\mathbb{F}}_{q}[T]$. In other words, $\overline{\phi}$ is a Drinfeld module of rank $R$ over ${\mathbb{F}}_{\mathfrak{p}}$.
It is immediate to see that if $\phi$ in the preceding definition is a Carlitz module, then it has good reduction everywhere. Next, we present the following theorem about a family of Drinfeld modules of Rank $1$.
@@
Let $P,h$ be polynomials in ${\mathbb{F}}_{q}[T]$ such that $P$ is monic irreducible with a residue field ${\mathbb{F}}_{\mathfrak{p}}$, and $g$ be a nonzero element in ${\mathbb{F}}_{q}(T)$ such that $v_{\mathfrak{p}}(g)=v_{\mathfrak{p}}(h)=0$. Denote by $\overline{g}\in{\mathbb{F}}_{\mathfrak{p}}$ the reduction of $g$ modulo $P$. Let $\phi^{(g)}:{\mathbb{F}}_{q}[T]\rightarrow{\mathbb{F}}_{q}(T)\{\tau\}$ be a Drinfeld module of rank $1$ where $\phi_{T}^{(g)}=T+g\tau$. Then, $P$ splits completely in the $h$-division field of $\phi$ if and only if $P\equiv\operatorname{Nr}_{{\mathbb{F}}_{\mathfrak{p}}/{\mathbb{F}}_{q}}(\overline{g})\mod h$.

### v2

@@ Abstract.
We introduce a new technique to construct rank-metric codes using the arithmetic theory of Drinfeld modules, and Dirichlet Theorem on polynomial arithmetic progressions. Using our methods, we obtain a new infinite family of optimal rank-metric codes with rank-locality, i.e. every code in our family achieves the information theoretical bound for rank-metric codes with rank-locality.
@@ 1. Introduction
Let $q$ be a prime power and $m$ a positive integer. Rank-metric codes are an important family of error correcting codes that are used for reliable communication over noisy channels [1, 5, 12, 18, 23] and for cryptography [7, 10, 21]. They have been widely studied in the literature due to their interesting algebraic combinatorial properties that allow to encode information in a way that the obtained codewords are resilient to rank errors, i.e. the noise is modeled as a sum of a low rank matrix.
@@
Rank-metric codes with rank-locality belong to an important subclass of rank-metric codes that allow efficient corrections of criss-cross failures in data centers (for additional detail on this see for example [14, Example 1]). Various constructions with flexible parameters are available for locally recoverable codes over the Hamming metric [4, 3, 6, 8, 11, 15, 16, 17] but only one for the rank-metric codes [14], which allows only very restrictive parameters (see for example the divisibility conditions in [14, Construction 1]). The difficulty that arises in the rank-metric context is that classical constructions fail due to both the non-commutative nature of the message space (which naturally consists of $q$-linearized polynomials with coefficients in ${\mathbb{F}}_{q^{m}}$), and the fact that one can control certain splitting conditions only for special linearized polynomials. For example, in [14, Construction 1], the kernels of the linearized polynomials used are orthogonal subspaces whose direct sum is a field, which gives restrictive divisibility conditions.
The purpose of this paper is to employ the theory of Drinfeld Modules to construct new rank metric codes with locality having parameters that are not available using other constructions. Our method employs a blend of supersingular Drinfeld Modules, Anderson motive decompositions, Dirichlet Theorem, and of course rank metric codes theory. Notice that differently from [22] (where the authors get classical locality, as explained in [14, Section 3]), we actually get proper rank-locality, as required in [14].
We structure the paper in three parts. The first one, Section 2, is merely introductory and provides the necessary background on Drinfeld Modules and Rank Metric Codes. The second one, consisting of Sections 3 and 4, provides the algebraic core of the construction. Then, in the third part we show that the hypothesis of our theorem can indeed be easily verified. In order to see this, we employ Dirichlet Theorem for polynomial arithmetic progressions to show the existence of an infinite family of our codes (Section 6), and provide two examples to show that the parameters can be kept small (Section 5).
@@ 2.1. Drinfeld Modules
In this subsection, we provide some background on the theory of Drinfeld modules. For a comprehensive treatise of the subject, the reader should refer to [19]. Throughout this section we assume $K\supseteq{\mathbb{F}}_{q}$ is a field and $a\in{\mathbb{F}}_{q}[T]$. Throughout the article, let $q$ be a prime power and ${\mathbb{F}}_{q}$ the finite field with $q$ elements. We denote by $A={\mathbb{F}}_{q}[T]$ the ring of polynomials over ${\mathbb{F}}_{q}$.
@@
Slightly abusing the notation, we also denote by $\mathfrak{p}$ the monic generator of $\mathfrak{p}=(P)$.
@@
Let $f=a_{h}\tau^{h}+a_{h+1}\tau^{h+1}\dots+a_{n}\tau^{n}\in K\{\tau\}$ such that $0\leq h\leq n$ with $a_{h}\neq 0,a_{n}\neq 0$. The height of $f$ is $\operatorname{ht}(f)=h$, and the $\tau$-degree (or simply the degree) of $f$ is $\deg(f)\coloneqq\deg_{\tau}(f)=n$. We also say that $f$ is separable if $\operatorname{ht}(f)=0$ and inseparable otherwise.
Definition 2.6.
@@
$a\longmapsto$ $\phi_{a}=\gamma(a)+\sum^{n}_{i=1}g_{i}(a)\tau^{i}$
where $g_{n}(a)\neq 0$ and $n=\operatorname{deg}_{T}(a)\cdot R$. The morphism $\gamma$ is called the structure morphism of $\phi$, and $\mathfrak{p}\coloneqq\ker(\gamma)$ is called the $A$-characteristic of $K$ and denoted by $\operatorname{char}_{A}(K)$.
For what follows, it is important to observe that $K$ could have different structures of $A$-field, namely it could be associated to different $A$-characteristics, depending on the choice of $\gamma$ and $\phi$. Observe also that for any $a=\sum_{i=0}^{n}a_{i}T^{i}$, we have $\phi_{a}=\sum_{i=0}^{n}a_{i}\phi_{T}^{i}$. In other words, $\phi_{T}$ determines $\phi$ entirely. Moreover, $K\{\tau\}$ is related to the ring of $q$-linearized polynomials $K\langle x\rangle$ via the following isomorphism
@@
$\iota(\phi_{a})$ is called the $a$-division polynomial of $\phi$ and denoted by $\phi_{a}(x)$.
Among all Drinfeld modules, there is a particular one that plays a special role in our paper. Therefore we introduce also the following.
Definition 2.8.
[19, Definition 3.2.5] The Carlitz module $C$ is the Drinfeld module of rank $1$ defined by $C_{T}=\gamma(T)+\tau$.
@@
Proposition 2.10.
[19, Lemma 3.2.11] If $\operatorname{char}_{A}(K)=\mathfrak{p}\neq 0$ and $\phi$ is a Drinfeld module of rank $R$ over $K$, then there is an integer $1\leq H(\phi)\leq R$, called the height of $\phi$, such that for all $0\neq a\in A$ we have
$\operatorname{ht}(\phi_{a})=H(\phi)\cdot\operatorname{ord}_{\mathfrak{p}}(a)\cdot\deg_{T}(\mathfrak{p})$
where $\operatorname{ord}_{\mathfrak{p}}(a)$ is the largest power of $\mathfrak{p}$ dividing $a$.
If a Drinfeld module $\phi$ of rank $R$ over a finite field field is such that $H(\phi)=R$, then $\phi$ is called supersingular. The interested reader should consult [19, Section 4.4] for further background and additional properties of supersingular Drinfeld modules. Such type of Drinfeld modules are quite special, as the following result shows.
Proposition 2.11.
Let $K\supseteq{\mathbb{F}}_{q}$ be a finite field, and $\phi:{\mathbb{F}}_{q}[T]\rightarrow K\{\tau\}$ be a Drinfeld module of rank $R$ and $\operatorname{char}_{A}(K)=\mathfrak{p}$. The followings are equivalent:
  1. (1)
    $\phi$ is supersingular (i.e. $H(\phi)=R)$
  2. (2)
    $\phi[\mathfrak{p}^{k}]=\{0\}$ for all $k\geq 1$
Proof.
Refer to [19, Theorem 4.4.1] for a proof and more equivalent conditions for a Drinfeld module to be supersingular. ∎
By definition of height, the simplest example of supersingular Drinfeld module is given by any Drinfeld module $\phi$ of rank 1, indeed $1\leq H(\phi)\leq R=1$. Another example is shown below.
Example 2.12.
[19, Example 4.4.5] Let $R$ be a prime, and $m$ be an integer coprime to $R$. Let $\alpha$ be a root of an irreducible polynomial $\mathfrak{p}\in{\mathbb{F}}_{q}[T]$ of degree $m$. The Drinfeld module $\phi:{\mathbb{F}}_{q}[T]\rightarrow{\mathbb{F}}_{q^{m}}\{\tau\}$ given by $\phi_{T}=\alpha+\tau^{R}$ is supersingular.
It turns out that the set of $a$-torsion points has a more general structure than just being an ${\mathbb{F}}_{q}$-vector space. The following result tells that $\phi[a]$ has an $A$-module structure, or equivalently that every $f\in\phi(A)$ acts as an endomorphism of $\phi[a]$.
@@
Proof.
Just notice that $\phi_{a}(\phi_{b}(\alpha))=\phi_{b}(\phi_{a}(\alpha))=\phi_{b}(0)=0$. ∎
@@
2.2. Rank Metric Codes
@@
One can check that this is indeed a distance on ${\mathbb{F}}^{m\times n}_{q}$.
@@
For any $x=(x_{1},\ldots,x_{n})\in{\mathbb{F}}^{n}_{q^{m}}$, we can define a matrix $M_{x}\in{\mathbb{F}}^{m\times n}_{q}$ associated with $x$ up to the choice of the basis as follows. Fix a basis $\mathcal{B}=\{b_{1},\ldots,b_{m}\}$ of ${\mathbb{F}}_{q^{m}}$ over ${\mathbb{F}}_{q}$. Then, $x_{j}=\sum_{i=1}^{m}c_{ij}b_{i}$ for some $c_{ij}\in{\mathbb{F}}_{q}$. Thus, $M_{x}=[c_{ij}]_{\begin{subarray}{c}1\leq i\leq m\\ 1\leq j\leq n\end{subarray}}$. In conclusion, we can define another type of rank-metric codes; namely (vector) rank-metric codes.
Definition 2.17.
Proposition 2.18.
$\dim_{{\mathbb{F}}_{q^{m}}}(\mathcal{C})\leq n-d_{R}(\mathcal{C})+1.$
Definition 2.19.
Definition 2.20.
Let $n,m,k$ be positive integers such that $m\geq n$. Let $\mathcal{P}=\{p_{1},\ldots,p_{n}\}$ be a set of elements in ${\mathbb{F}}_{q^{m}}$ that are ${\mathbb{F}}_{q}$-linearly independent. Let $\bm{m}=(\bm{m}_{0},\ldots,\bm{m}_{k-1})\in{\mathbb{F}}^{k}_{q^{m}}$ and $\{G_{\bm{m}}(x)=\sum_{j=0}^{k-1}\bm{m}_{j}x^{[j]}\}$ be the set of linearized polynomials of $q$-degree at most k1k-1 in ${\mathbb{F}}_{q^{m}}[x]$. We define the encoding map as
$\bm{m}\longmapsto$ $\{G_{\bm{m}}(\gamma),\gamma\in\mathcal{P}\}.$
Definition 2.21.
  • $i\in\Gamma(i)$
  • $|\Gamma(i)|\leq r+\delta-1$
  • $d_{R}(\mathcal{C}\mid_{\Gamma(i)})\geq\delta$
Proposition 2.22.
Proof.
See [14, Theorem 1]
3. General Theorem and Main Construction
For this section and the theorems within, we will use the following notation so we will be able to make the statements of each theorem lighter. Let $q$ be a prime power, and ${\mathbb{F}}_{q}$ be the finite field with $q$ elements. For $m,t>0$ let ${\mathbb{F}}_{q^{m}}$ be the algebraic field extension of ${\mathbb{F}}_{q}$ of degree $m$, and $K$ be a finite extension of ${\mathbb{F}}_{q^{m}}$ of degree $t$ (namely $K:{\mathbb{F}}_{q}]=tm)$. We denote with $A\coloneqq{\mathbb{F}}_{q}[T]$ the ring of polynomials in $T$ with coefficient over ${\mathbb{F}}_{q}$. Let $0<\ell<q$, $r\geq 1$ and $\delta\geq 2$ be positive integers, and define $R\coloneqq r+\delta-1$.
@@ 3. General Theorem and Main Construction
Choose distinct $a_{1},\ldots,a_{\ell}\in{\mathbb{F}}_{q}^{\times}$ and define the polynomial $h=\prod_{i=1}^{\ell}(T-a_{i})\in A$. Let $s$ be such that $s+1\leq\ell$. Let $\phi:A\rightarrow{\mathbb{F}}_{q^{m}}\{\tau\}$ be a Drinfeld module of rank $R$, with $A$-characteristic being a monic irreducible polynomial $\mathfrak{p}$ of degree $m$ such that $\gcd(\mathfrak{p},h)=1$. Assume ${\mathbb{F}}_{q^{m}}(\phi[h])\subseteq K$. Then, there exists a
$\left(tm\times\ell R,tm(s+1)r,\ell R-Rs-r+1,r,\delta\right)$
(matrix) rank-metric code that is optimal w.r.t. the bound shown in Proposition 2.22.
In subsection 3.1 we will prove Theorem 3.1. Knowing that ${\mathbb{F}}_{q^{m}}(\phi[h])\subseteq K$ is fundamental to construct our code. In Section 4 we will show the existence of at least two families of Drinfeld modules for which such assumption is always verified. This will amount to finding an irreducible polynomial within certain wanted residue class modulo $h$. In general this is possible even for small $m$ (see Section 5.1 and 5.2 for examples). In Section 6 we will use Dirichlet Theorem for polynomial arithmetic progressions to provide a proof that there exists a wide range of parameters for which we can guarantee the existence of a monic irreducible $\mathfrak{p}$ such that ${\mathbb{F}}_{q^{m}}(\phi[h])\subseteq K$.
Lemma 3.2.
With notation and hypothesis as in Theorem 3.1, we have that
$\dim_{{\mathbb{F}}_{q}}\phi[h]=\ell R.$
In particular $\dim_{{\mathbb{F}}_{q}}\phi[T-a_{i}]=R.$